# Multi-fidelity Bayesian Optimization: A Review

### v2

@@ 1 Introduction
The mathematical modeling of problem (1) utilizes a mathematical model derived from our decent understanding of the governing equations of the physical process underlying the relationship between the design objective and design variables; see Fig. 1. The expression of this mathematical model is the objective function $f(\bf x)$. Unfortunately, solving the governing equations of $f(\bf x)$ is often analytically intractable in engineering design. Thus, experiments are considered as the ground truth to measure certain aspects of the physical process. However, they are subject to uncertainties in design manufacturing, operational control, and measurement errors, which must be considered in experimental data analysis. Oftentimes the mathematical model is implemented via its computational counterparts, such as finite element (FE) models [4], which simulate the physical process under hypothetical scenarios.
Depending on their prediction qualities and costs, computational models for the physical process can be classified into high- and low-fidelity models; see Fig. 1. A high-fidelity model (HF model) is a computational model that predicts the performance metrics with an accuracy level sufficient for a basic understanding of the underlying physical process and evaluation of the objective function. We denote $f_{\text{H}}(\bf x)$ as the objective function value at ${\bf x}$ obtained from experiments or an HF model. We consider both data from experiments and evaluations of an HF model as HF data. Meanwhile, a low-fidelity model (LF model) is a computational model that is designated for applications where a lower level of accuracy is acceptable for evaluating the objective function. The objective function and data from predictions of an LF model are $f_{\text{L}}(\bf x)$ and LF data, respectively. The advantage of the LF model is that it can be much more computationally efficient than its HF counterpart, both in space and time costs. Examples of LF models include simplified-physics approximations and projection-based reduced order models (ROMs), as described in [5].
@@
Once the mathematical modeling of the objective function is settled, we can consider the selection of an optimization algorithm. Ideally, a good optimization algorithm should possess three properties: robustness, accuracy, and efficiency [6]. However, these goals may conflict and trade-offs between them should be addressed. A prudent approach is to select optimization algorithms based on how the objective function is accessed; see Fig. 2.
@@
As a successful approach to addressing the efficiency-accuracy trade-off, surrogate-based optimization (SbO) approximates the objective function $f(\bf x)$ by a surrogate model [9, 7, 10, 11, 12]. The surrogate can be a data-fit model constructed by fitting the parameters of an interpolation, regression, or machine learning model, given pairs of design variables and performance metrics or the overall score derived from HF data, thereby favoring the accuracy aspect. Alternatively, it can be induced from an LF model used as an approximation of the physical process in Fig. 1, which, however, favors the efficiency aspect. We can further classify surrogates into deterministic surrogates, such as response surfaces and most neural networks, and probabilistic surrogates, such as Gaussian process (GP) [13]. We can also categorize them into local and global surrogates. A local surrogate approximates a black-box objective function in a restricted region of the design variable space, e.g., the neighborhood of a design variable vector. A global surrogate approximates all predictions of that black-box objective function.
@@
Bayesian optimization (BO) is a global optimization technique that assumes a prior probabilistic model on the objective function, and combines it with the available data to guide the optimization process. BO is well suited for small- to moderate-dimensional optimization problems that involve expensive-to-evaluate objective functions that offer no efficient mechanism to estimate their derivatives [24, 25, 26]. Most of the time, a GP is used as the prior model in BO, due to its tractability and flexibility. Typically, the GP prior is specified simply by a constant mean function and a covariance function with a closed form and a few hyperparameters. Given the dataset at each optimization iteration, the posterior model (i.e., the prior model conditioned on the dataset) determines an acquisition function reflecting an optimization policy under uncertainty, which is then numerically maximized to find the next design point. A derivative-free, global optimization algorithm is usually used for this purpose. The GP posterior at each iteration serves as a surrogate for the objective function, and therefore BO is a type of SbO. BO tends to favor accuracy if it constructs the GP posterior based on HF data. The details of BO and its applications are deferred until Section 3.2.
@@
We organize the rest of this paper as follows. Section 2 surveys two basic elements of MFO: MF modeling methods and MFO strategies. Section 3 provides an overview of MF BO. Section 4 describes important GP-based MF surrogates. Section 5 surveys various acquisition functions of BO and how we can modify them for use of MF BO. Section 6 shows the applications of several MF BO methods for optimizing benchmark functions and standard airfoil shapes. Section 7 reviews the extensions of BO and MFO to address intricate yet important design optimization problems, including constrained optimization, high-dimensional optimization, optimization under uncertainty, and multi-objective optimization. Section 8 summarizes and concludes this paper.
@@ 2 Multi-fidelity optimization
This section starts with a brief overview of SbO (Section 2.1). It then surveys two elements of MFO: MF surrogates (Section 2.2) and MFO strategies (Section 2.3).
@@ 2.1 Surrogate-based optimization
The use of MFO in engineering design was initiated by Haftka [43] with a multiplicative MF model (Section 2.2). This model broadens the application range of HF responses by adjusting the associated LF responses. During the optimization process, the multiplicative MF model is imposed consistency conditions, which require that the HF response and its derivatives at a given design point must align with those of the corresponding scaled LF response [44]. If the HF derivatives are not available, the consistency can be defined for the response only [45].
Early attempts at MFO relied on local approximations with deterministic MF surrogates valid in the neighborhood of the current point of design variables. Applications of local MFO methods have been found in supersonic transport design [46], aerodynamic wing design [44], and airfoil design [47]. Since the seminal works by Huang et al. [15] and Forrester et al. [16], MFO has shifted its attention to global approximations with probabilistic MF surrogates valid for a large region of the design variable space. As we will see in Section 2.3, a majority of recent applications of MFO in engineering design have been from global methods.
@@ 2.2 Multi-fidelity surrogates
Note that the form in Eq. 2 does not include cases when multiple LF models are available. However, we can use it to recover almost all MF surrogates derived from multiple hierarchical LF models listed in a recent review of MF modeling methods [48].
@@
Based on the values of $w$ and $c$, we can further classify adjustment MF surrogates into three types as shown in Fig. 3:
@@
Table 1 enumerates important published works for each type of adjustment MF surrogates. We see that, a substantial portion of these works has relied on additive adjustment MF surrogates.
@@
Apart from the form in Eq. 2, adjustment MF surrogates can also be described using weighted average models when there exist a total of $m$ LF models, such that [41]
@@
where $f_{i}({\bf x})$ represents a surrogate for approximating the prediction of LF model $i$, $w_{i}$ is a weight value corresponding to $f_{i}({\bf x})$ with $\sum_{i=1}^{m}w_{i}=1$, and $\rho_{i}({\bf x})$ is a weight function corresponding to $f_{i}({\bf x})$ with $\sum_{i=1}^{m}\rho_{i}({\bf x})=1$. We use Eq. 3a when the weights are fixed in different regions of the design variable space. We use Eq. 3b when there is a correlation between $f_{\text{H}}({\bf x})$ and $f_{i}({\bf x})$ at any values of the design variables.
Based on how $\rho({\bf{x}})$ and $\delta({\bf{x}})$ are modeled, we can further classify adjustment MF surrogates into deterministic and probabilistic ones; see Fig. 4. Table 2 lists notable published works for each of these classes.
@@
Deterministic adjustment MF surrogates described in Fig. 4 model both $\rho({\bf{x}})$ and $\delta({\bf{x}})$ as linear combinations of a finite number of basis functions $\zeta_{i}({\bf{x}})$ and $\xi_{j}({\bf x})$, respectively. Popular basis functions include monomials [53, 54], neural networks [56], radial basis functions [89, 103, 58], and orthogonal polynomial functions [52]. The combination coefficients and hyperparameters associated with the basis functions are determined by the least squares or regularized least squares approach. In comparison, probabilistic adjustment MF surrogates describe $\rho({\bf{x}})$ or $\delta({\bf{x}})$ as a stochastic process, which is often a GP characterized by a mean function and a covariance function. Several GP-based MF surrogates can be found in Brevault et al. [107]. The mathematical expressions of common GP-based MF surrogates are deferred until Section 4.
@@
[…]
@@
The class of input-augmentation MF modeling methods describes MF surrogates as functions that depend not only on design variables but also on fidelity variables. This differs from the class of adjustment MF surrogates in Eqs. (2) and (3) which are functions of the design variables only. The fidelity variables are indeed categorical but can be assumed to be continuous for computational efficiency [117]. If categorical fidelity variables are considered, non-continuous covariance functions are proposed for constructing GP-based MF surrogates. Section 4.6 provides different input-augmentation MF modeling methods for both continuous and categorical fidelity variables. Note that most input-augmentation MF surrogates in the literature are probabilistic and have been predominantly used in BO.
@@ 2.3 Multi-fidelity optimization strategies
MFO strategies take advantage of MF surrogates and/or their derivatives to accelerate the optimization process. Each MFO strategy is characterized by its way of updating the design points and MF surrogates during the optimization process. There are three approaches to MFO strategies: derivative-based, model-then-optimize, and model-and-optimize. Notable published works on the algorithms of each approach are listed in Table 4.
@@
This approach, also known as sequential model-based optimization approach [166], iterates between updating global MF surrogates for the costly objective function and using them to propose new design points via maximizing an acquisition function (i.e., infill criterion or figure of merit). From this definition, MF BO is a type of model-and-optimization approach. The model-and-optimize approach, via formulating the acquisition function, minimizes the need for using numerous optimization parameters as the derivative-based approach does. Factors that affect the performance of model-and-optimize algorithms include the choice of MF surrogates, the selection of acquisition functions, and the construction of initial MF surrogates. As shown in Table 4, this approach has recently attracted a large number of published works on the applications of MFO in design optimization.
@@ 3 Multi-fidelity Bayesian optimization
As illustrated in Fig. 2 and briefly defined in Section 1, BO is a global optimization technique working based on constructing a probabilistic surrogate model to represent our belief about an expensive-to-evaluate objective function given its observations, which then defines an optimization policy via an acquisition function that is maximized before a new design point is selected for evaluating the objective function to update our belief. BO was originated by Močkus [167] and popularized by Jones et al. [168] and their work on the Efficient Global Optimization (EGO) algorithm. Combining BO with MFO leads to MF BO which further accelerates the optimization process.
This section starts with describing the fundamental elements of EGO (Section 3.1) and BO (Section 3.2), and how the two concepts connect to each other. It then elucidates distinctive characteristics that set MF BO apart from BO (Section 3.3).
@@ 3.1 Efficient global optimization
In the second step, EGO formulates what is known as the expected improvement based on three sources of information: (1) the best design point found in the current training dataset, (2) the best linear unbiased predictor, and (3) the MSE of the predictor. The concept of expected improvement (Section 5.1.1) provides a delicate balance between exploiting the information provided by the predictor and exploring uncertain regions within it. Specifically, exploitation (i.e., immediate reward) focuses on regions of the design variable space where the objective function values are expected to be small, while exploration (i.e., expected future reward) discovers regions where the predictions of the objective function are highly uncertain.
@@ 3.2 Bayesian optimization
1:Input: $f_{\text{H}}(\cdot)$, $\mathcal{X}$, $K$, $N$;
2:Generate $N$ samples of ${\bf x}$;
@@
Figure 5 illustrates two consecutive iterations of BO for minimizing a univariate objective function $f(x)$. In Fig. 5(a), BO fits a GP model for $f(x)$ to a training dataset of six samples (top panel). Subsequently, it formulates and maximizes an acquisition function $\alpha(x)$ to identify a new design point (bottom panel). In Fig. 5(b), BO proceeds by updating the GP model and reformulating $\alpha(x)$ with the design point identified in the previous iteration.
@@
Algorithm 1 provides a pseudo-code for solving problem (1) using the generic BO, where the objective function is the prediction $f_{\text{H}}(\bf x)$ of an HF model. Parameters $K$ and $N$ in 1 represent threshold values for the number of BO iterations and the number of initial training samples, respectively. $\mathcal{D}^{0}$ in 6 is the initial training dataset generated from the HF model. Note that $\mathcal{D}^{0}$ can be HF data, which contain both predictions of the HF model and measures from experiments. $\hat{f}_{\text{H}}^{k}({\bf x})$ in 9 and $\mathcal{D}^{k}$ in 13 are the GP posterior and training dataset associated with iteration $k$ of BO, respectively. The acquisition function $\alpha({\bf x})$ is formulated in 10 based on $\mathcal{D}^{k-1}$ and $\hat{f}_{\text{H}}^{k}({\bf x})$, which is detailed in Section 5. Maximizing $\alpha({\bf x})$ in 11 is often straightforward as it does not invoke the HF model. By doing so, we replace problem (1), which is difficult to solve, with a series of simpler, inexpensive problems of maximizing the acquisition function in 11.
@@
At this point, a critical question arises: under what conditions can we ensure that the existence of a unique global solution to a GP-distributed objective function ${f}_{\text{H}}({\bf x})$? Addressing this question is crucial because it makes the construction of the GP model $\hat{f}_{\text{H}}^{k}({\bf x})$ in 9 of Algorithm 1 meaningful. Since the optimization problem is completely defined by the mean and covariance functions of the GP and the design domain $\mathcal{X}$, their properties, intuitively, decide whether a unique global solution is guaranteed. In fact, we can be certain that a global solution exists and that is unique under the following two mild assumptions: (1) the mean and covariance functions are continuous over a compact design domain $\mathcal{X}$, and (2) there are no two unique points in the domain that can have perfectly correlated function values [173]. The reader may consult Section 2.7 of Garnett [173] for detailed discussions on the existence and uniqueness of global solutions to GP-distributed functions.
@@
The unique ability to optimize expensive-to-evaluate objective functions, i.e., making efficient use of the available data for ensuring solution accuracy, has made BO an invaluable tool for various applications in scientific and engineering design. These applications have been found in machine learning [174], aircraft design [175], material design [176, 30], experimental design [177], material science [178], structural engineering [179], transportation [180], chemical engineering [181], electronics engineering [182], environmental engineering [183], and physics [184], to name a few. Moreover, sophisticated BO algorithms have been developed to (1) examine the performance of acquisition functions formulated from different optimization policies when optimizing objective functions in the face of uncertainty and (2) extend applications of BO to a wide range of design optimization problems. The first aspect has examined the performance of BO acquisition functions derived from different formulation approaches: improvement-based [185], optimistic [186], information-based  [187], and likelihood-weighted [188]. It has also explored the performance of acquisition functions from one-step look-ahead and multi-step look-ahead perspectives (Section 5.1). The second aspect has focused on developing advanced BO algorithms to solve intricate yet important design optimization problems such as constrained problems, high-dimensional problems, problems under uncertainty, and multi-objective problems (Section 7). These problems are difficult to solve because of their own nature and/or limitations of the generic BO. For the details of recent advances in BO, the reader may refer to comprehensive surveys and tutorials by Brochu et al. [189], Shahriari et al. [25], Frazier [26], and Wang et al. [31].

### v3

@@ 1 Introduction
The mathematical modeling of problem (1) utilizes a mathematical model derived from our decent understanding of the governing equations of the physical process underlying the relationship between the design objective and design variables; see Fig.˜1. The expression of this mathematical model is the objective function $f(\bf x)$. Unfortunately, solving the governing equations of $f(\bf x)$ is often analytically intractable in engineering design. Thus, experiments are considered as the ground truth to measure certain aspects of the physical process. However, they are subject to uncertainties in design manufacturing, operational control, and measurement errors, which must be considered in experimental data analysis. Oftentimes the mathematical model is implemented via its computational counterparts, such as finite element (FE) models [4], which simulate the physical process under hypothetical scenarios.
Depending on their prediction qualities and costs, computational models for the physical process can be classified into high- and low-fidelity models; see Fig.˜1. A high-fidelity model (HF model) is a computational model that predicts the performance metrics with an accuracy level sufficient for a basic understanding of the underlying physical process and evaluation of the objective function. We denote $f_{\text{H}}(\bf x)$ as the objective function value at $\bf x$} obtained from experiments or an HF model. We consider both data from experiments and evaluations of an HF model as HF data. Meanwhile, a low-fidelity model (LF model) is a computational model that is designated for applications where a lower level of accuracy is acceptable for evaluating the objective function. The objective function and data from predictions of an LF model are $f_{\text{L}}(\bf x)$ and LF data, respectively. The advantage of the LF model is that it can be much more computationally efficient than its HF counterpart, both in space and time costs. Examples of LF models include simplified-physics approximations and projection-based reduced order models (ROMs), as described in [5].
@@
Once the mathematical modeling of the objective function is settled, we can consider the selection of an optimization algorithm. Ideally, a good optimization algorithm should possess three properties: robustness, accuracy, and efficiency [6]. However, these goals may conflict and trade-offs between them should be addressed. A prudent approach is to select optimization algorithms based on how the objective function is accessed; see Fig.˜2.
@@
As a successful approach to addressing the efficiency-accuracy trade-off, surrogate-based optimization (SbO) approximates the objective function $f(\bf x)$ by a surrogate model [9, 7, 10, 11, 12]. The surrogate can be a data-fit model constructed by fitting the parameters of an interpolation, regression, or machine learning model, given pairs of design variables and performance metrics or the overall score derived from HF data, thereby favoring the accuracy aspect. Alternatively, it can be induced from an LF model used as an approximation of the physical process in Fig.˜1, which, however, favors the efficiency aspect. We can further classify surrogates into deterministic surrogates, such as response surfaces and most neural networks, and probabilistic surrogates, such as Gaussian process (GP) [13]. We can also categorize them into local and global surrogates. A local surrogate approximates a black-box objective function in a restricted region of the design variable space, e.g., the neighborhood of a design variable vector. A global surrogate approximates all predictions of that black-box objective function.
@@
Bayesian optimization (BO) is a global optimization technique that assumes a prior probabilistic model on the objective function, and combines it with the available data to guide the optimization process. BO is well suited for small- to moderate-dimensional optimization problems that involve expensive-to-evaluate objective functions that offer no efficient mechanism to estimate their derivatives [24, 25, 26]. Most of the time, a GP is used as the prior model in BO, due to its tractability and flexibility. Typically, the GP prior is specified simply by a constant mean function and a covariance function with a closed form and a few hyperparameters. Given the dataset at each optimization iteration, the posterior model (i.e., the prior model conditioned on the dataset) determines an acquisition function reflecting an optimization policy under uncertainty, which is then numerically maximized to find the next design point. A derivative-free, global optimization algorithm is usually used for this purpose. The GP posterior at each iteration serves as a surrogate for the objective function, and therefore BO is a type of SbO. BO tends to favor accuracy if it constructs the GP posterior based on HF data. The details of BO and its applications are deferred until Section˜3.2.
@@
We organize the rest of this paper as follows. Section˜2 surveys two basic elements of MFO: MF modeling methods and MFO strategies. Section˜3 provides an overview of MF BO. Section˜4 describes important GP-based MF surrogates. Section˜5 surveys various acquisition functions of BO and how we can modify them for use of MF BO. Section˜6 shows the applications of several MF BO methods for optimizing benchmark functions and standard airfoil shapes. Section˜7 reviews the extensions of BO and MFO to address intricate yet important design optimization problems, including constrained optimization, high-dimensional optimization, optimization under uncertainty, and multi-objective optimization. Section˜8 summarizes and concludes this paper.
@@ 2 Multi-fidelity optimization
This section starts with a brief overview of SbO (Section˜2.1). It then surveys two elements of MFO: MF surrogates (Section˜2.2) and MFO strategies (Section˜2.3).
@@ 2.1 Surrogate-based optimization
The use of MFO in engineering design was initiated by Haftka [43] with a multiplicative MF model (Section˜2.2). This model broadens the application range of HF responses by adjusting the associated LF responses. During the optimization process, the multiplicative MF model is imposed consistency conditions, which require that the HF response and its derivatives at a given design point must align with those of the corresponding scaled LF response [44]. If the HF derivatives are not available, the consistency can be defined for the response only [45].
Early attempts at MFO relied on local approximations with deterministic MF surrogates valid in the neighborhood of the current point of design variables. Applications of local MFO methods have been found in supersonic transport design [46], aerodynamic wing design [44], and airfoil design [47]. Since the seminal works by Huang et al. [15] and Forrester et al. [16], MFO has shifted its attention to global approximations with probabilistic MF surrogates valid for a large region of the design variable space. As we will see in Section˜2.3, a majority of recent applications of MFO in engineering design have been from global methods.
@@ 2.2 Multi-fidelity surrogates
Note that the form in Eq.˜2 does not include cases when multiple LF models are available. However, we can use it to recover almost all MF surrogates derived from multiple hierarchical LF models listed in a recent review of MF modeling methods [48].
@@
Based on the values of $w$ and $c$, we can further classify adjustment MF surrogates into three types as shown in Fig.˜3:
@@
Table˜1 enumerates important published works for each type of adjustment MF surrogates. We see that, a substantial portion of these works has relied on additive adjustment MF surrogates.
@@
Apart from the form in Eq.˜2, adjustment MF surrogates can also be described using weighted average models when there exist a total of $m$ LF models, such that [41]
@@
where $f_{i}({\bf x})$ represents a surrogate for approximating the prediction of LF model $i$, $w_{i}$ is a weight value corresponding to $f_{i}({\bf x})$ with $\sum_{i=1}^{m}w_{i}=1$, and $\rho_{i}({\bf x})$ is a weight function corresponding to $f_{i}({\bf x})$ with $\sum_{i=1}^{m}\rho_{i}({\bf x})=1$. We use Eq.˜3a when the weights are fixed in different regions of the design variable space. We use Eq.˜3b when there is a correlation between $f_{\text{H}}({\bf x})$ and $f_{i}({\bf x})$ at any values of the design variables.
Based on how $\rho({\bf{x}})$ and $\delta({\bf{x}})$ are modeled, we can further classify adjustment MF surrogates into deterministic and probabilistic ones; see Fig.˜4. Table˜2 lists notable published works for each of these classes.
@@
Deterministic adjustment MF surrogates described in Fig.˜4 model both $\rho({\bf{x}})$ and $\delta({\bf{x}})$ as linear combinations of a finite number of basis functions $\zeta_{i}({\bf{x}})$ and $\xi_{j}({\bf x})$, respectively. Popular basis functions include monomials [53, 54], neural networks [56], radial basis functions [89, 103, 58], and orthogonal polynomial functions [52]. The combination coefficients and hyperparameters associated with the basis functions are determined by the least squares or regularized least squares approach. In comparison, probabilistic adjustment MF surrogates describe $\rho({\bf{x}})$ or $\delta({\bf{x}})$ as a stochastic process, which is often a GP characterized by a mean function and a covariance function. Several GP-based MF surrogates can be found in Brevault et al. [107]. The mathematical expressions of common GP-based MF surrogates are deferred until Section˜4.
@@
The class of composition methods consists of input-input and output-output mapping MF techniques. MF surrogates from input-input mapping techniques often take the form of $f_{\text{H}}({\bf x})\approx f_{\text{L}}(g({\bf x}))$, where $g(\cdot)$ maps the input space of the HF model to that of the LF model. Given an initial set of HF design variable values and the corresponding output values, the set of corresponding LF design variable values is found by a parameter extraction, and a deterministic mapping $g(\cdot)$ can be found iteratively by defining $g^{j}(\cdot)$ at iteration $j$ as a linear combination of some predefined and fixed basis functions [21]. An appropriate mapping $g(\cdot)$ is found when $\|{\bf f}_{\text{H}}(\cdot)-{\bf f}_{\text{L}}(\cdot)\|\leq\varepsilon$, where ${\bf f}_{\text{L}}$ corresponds to $g$ values evaluated at the given HF design variable points, $\|\cdot\|$ the $L_{2}$ norm, and $\varepsilon$ a small positive constant . Meanwhile, MF surrogates from output-output mapping techniques take the form of $f_{\text{H}}({\bf x})\approx h\left(f_{\text{L}}({\bf x})\right)$, where $h(\cdot)$ maps the output space of the LF model to that of the HF model. The idea is to convert the original many-to-one mapping relationship between the space of HF design variables and the space of HF objective function to a one-to-one mapping relationship between certain regions of the spaces of LF and HF objective functions. This is possible because the values of LF objective function $f_{\text{L}}(\cdot)$ at HF design variable values can be obtained easily at a low computational cost. Moreover, a composition MF surrogate can be deterministic or probabilistic depending on how we define the mappings $g(\cdot)$ and $h(\cdot)$. The reader is referred to Table˜3 for a list of published works related to the composition MF modeling methods.
@@
The class of input-augmentation MF modeling methods describes MF surrogates as functions that depend not only on design variables but also on fidelity variables. This differs from the class of adjustment MF surrogates in Eqs. (2) and (3) which are functions of the design variables only. The fidelity variables are indeed categorical but can be assumed to be continuous for computational efficiency [117]. If categorical fidelity variables are considered, non-continuous covariance functions are proposed for constructing GP-based MF surrogates. Section˜4.6 provides different input-augmentation MF modeling methods for both continuous and categorical fidelity variables. Note that most input-augmentation MF surrogates in the literature are probabilistic and have been predominantly used in BO.
@@ 2.3 Multi-fidelity optimization strategies
MFO strategies take advantage of MF surrogates and/or their derivatives to accelerate the optimization process. Each MFO strategy is characterized by its way of updating the design points and MF surrogates during the optimization process. There are three approaches to MFO strategies: derivative-based, model-then-optimize, and model-and-optimize. Notable published works on the algorithms of each approach are listed in Table˜4.
@@
This approach, also known as sequential model-based optimization approach [166], iterates between updating global MF surrogates for the costly objective function and using them to propose new design points via maximizing an acquisition function (i.e., infill criterion or figure of merit). From this definition, MF BO is a type of model-and-optimization approach. The model-and-optimize approach, via formulating the acquisition function, minimizes the need for using numerous optimization parameters as the derivative-based approach does. Factors that affect the performance of model-and-optimize algorithms include the choice of MF surrogates, the selection of acquisition functions, and the construction of initial MF surrogates. As shown in Table˜4, this approach has recently attracted a large number of published works on the applications of MFO in design optimization.
@@ 3 Multi-fidelity Bayesian optimization
As illustrated in Fig.˜2 and briefly defined in Section˜1, BO is a global optimization technique working based on constructing a probabilistic surrogate model to represent our belief about an expensive-to-evaluate objective function given its observations, which then defines an optimization policy via an acquisition function that is maximized before a new design point is selected for evaluating the objective function to update our belief. BO was originated by Močkus [167] and popularized by Jones et al. [168] and their work on the Efficient Global Optimization (EGO) algorithm. Combining BO with MFO leads to MF BO which further accelerates the optimization process.
This section starts with describing the fundamental elements of EGO (Section˜3.1) and BO (Section˜3.2), and how the two concepts connect to each other. It then elucidates distinctive characteristics that set MF BO apart from BO (Section˜3.3).
@@ 3.1 Efficient global optimization
In the second step, EGO formulates what is known as the expected improvement based on three sources of information: (1) the best design point found in the current training dataset, (2) the best linear unbiased predictor, and (3) the MSE of the predictor. The concept of expected improvement (Section˜5.1.1) provides a delicate balance between exploiting the information provided by the predictor and exploring uncertain regions within it. Specifically, exploitation (i.e., immediate reward) focuses on regions of the design variable space where the objective function values are expected to be small, while exploration (i.e., expected future reward) discovers regions where the predictions of the objective function are highly uncertain.
@@ 3.2 Bayesian optimization
1:Input: $f_{\text{H}}(\cdot)$, $\mathcal{X}$, $K$, NN;
2:Generate NN samples of $\bf x$};
@@
Figure˜5 illustrates two consecutive iterations of BO for minimizing a univariate objective function $f(x)$. In Fig.˜5(a), BO fits a GP model for $f(x)$ to a training dataset of six samples (top panel). Subsequently, it formulates and maximizes an acquisition function $\alpha(x)$ to identify a new design point (bottom panel). In Fig.˜5(b), BO proceeds by updating the GP model and reformulating $\alpha(x)$ with the design point identified in the previous iteration.
@@
Algorithm˜1 provides a pseudo-code for solving problem (1) using the generic BO, where the objective function is the prediction $f_{\text{H}}(\bf x)$ of an HF model. Parameters $K$ and $N$ in Line 1 represent threshold values for the number of BO iterations and the number of initial training samples, respectively. $\mathcal{D}^{0}$ in Line 6 is the initial training dataset generated from the HF model. Note that $\mathcal{D}^{0}$ can be HF data, which contain both predictions of the HF model and measures from experiments. $\hat{f}_{\text{H}}^{k}({\bf x})$ in Line 9 and $\mathcal{D}^{k}$ in Line 13 are the GP posterior and training dataset associated with iteration $k$ of BO, respectively. The acquisition function $\alpha({\bf x})$ is formulated in Line 10 based on $\mathcal{D}^{k-1}$ and $\hat{f}_{\text{H}}^{k}({\bf x})$, which is detailed in Section˜5. Maximizing $\alpha({\bf x})$ in Line 11 is often straightforward as it does not invoke the HF model. By doing so, we replace problem (1), which is difficult to solve, with a series of simpler, inexpensive problems of maximizing the acquisition function in Line 11.
@@
At this point, a critical question arises: under what conditions can we ensure that the existence of a unique global solution to a GP-distributed objective function ${f}_{\text{H}}({\bf x})$? Addressing this question is crucial because it makes the construction of the GP model $\hat{f}_{\text{H}}^{k}({\bf x})$ in Line 9 of Algorithm˜1 meaningful. Since the optimization problem is completely defined by the mean and covariance functions of the GP and the design domain $\mathcal{X}$, their properties, intuitively, decide whether a unique global solution is guaranteed. In fact, we can be certain that a global solution exists and that is unique under the following two mild assumptions: (1) the mean and covariance functions are continuous over a compact design domain $\mathcal{X}$, and (2) there are no two unique points in the domain that can have perfectly correlated function values [173]. The reader may consult Section 2.7 of Garnett [173] for detailed discussions on the existence and uniqueness of global solutions to GP-distributed functions.
@@
The unique ability to optimize expensive-to-evaluate objective functions, i.e., making efficient use of the available data for ensuring solution accuracy, has made BO an invaluable tool for various applications in scientific and engineering design. These applications have been found in machine learning [174], aircraft design [175], material design [176, 30], experimental design [177], material science [178], structural engineering [179], transportation [180], chemical engineering [181], electronics engineering [182], environmental engineering [183], and physics [184], to name a few. Moreover, sophisticated BO algorithms have been developed to (1) examine the performance of acquisition functions formulated from different optimization policies when optimizing objective functions in the face of uncertainty and (2) extend applications of BO to a wide range of design optimization problems. The first aspect has examined the performance of BO acquisition functions derived from different formulation approaches: improvement-based [185], optimistic [186], information-based  [187], and likelihood-weighted [188]. It has also explored the performance of acquisition functions from one-step look-ahead and multi-step look-ahead perspectives (Section˜5.1). The second aspect has focused on developing advanced BO algorithms to solve intricate yet important design optimization problems such as constrained problems, high-dimensional problems, problems under uncertainty, and multi-objective problems (Section˜7). These problems are difficult to solve because of their own nature and/or limitations of the generic BO. For the details of recent advances in BO, the reader may refer to comprehensive surveys and tutorials by Brochu et al. [189], Shahriari et al. [25], Frazier [26], and Wang et al. [31].